karleigh:

Rachel works at a bookstore. On Tuesday, she sold twice as many books as she did on Monday. On Wednesday, she sold 6 fewer books than she did on Tuesday. During the three days Rachel sold 19 books. Create an equation that can be used to find m, the number of books Rachel sold on Monday.

darkknight:

let m=monday, t=tuesday and W=wednesday So T = 2m W = T-6 M+T+w = 19 So we can substitute 2m into the equation above for T, using substitution, So that M + 2m + w = 19 3m + w = 19 We also have W = T-6, And we can plug in 2m for T So W = (2m)-6 Now if we use substitution again, Than 3m+2m-6 = 19 5m = 25 m=5 That is the process for getting the books sold on Monday, lines 2-4 are the equations. Hope this helps and have a good day :), also be sure to click the best response button but that Is entirely up to you.

YNWMelly:

...to be continued